OpenStudy (anonymous):

the base of a triangle exceeds the height by 5 inches. if the area is 75 square inches, find the length and the base of the height of the traingle

OpenStudy (campbell_st):

well let the height by h and the base is h + 5  the formula for the area is \[75 = \frac{1}{2} \times (h + 5) \times h\] just solve for h

OpenStudy (anonymous):

is it height 10 and length of base 15?

OpenStudy (campbell_st):

seems to fit the conditions nicely, well done